# Positive Ricci Curvature on Twisted Suspensions

### v2

@@ 1. Introduction and Main Results
$G\cong{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k_{1_{\phantom{1}}}\mathbb{Z}$}\right.}\oplus\dots\oplus{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}\right.}\oplus\mathbb{Z}^{\ell_{2}}$
@@ Acknowledgements.
The author would like to thank Fernando Galaz-García and Philippe Kupper for helpful discussions, and them as well as Sam Hagh Shenas Noshari and Michael Wiemeler for helpful comments on an earlier version of this article. The author would also like to thank the Department of Mathematical Sciences of Durham University for its hospitality during a stay where parts of this work were carried out.
@@ 2.2. Twisted suspensions
In [16] another suspension operation, denoted $\widetilde{\Sigma}_{e}$, is introduced, which is obtained by replacing the identity map on $S^{n-1}\times S^{1}$ in Definition 2.4 by a map $\tilde{\alpha}\colon S^{n-1}\times S^{1}\to S^{n-1}\times S^{1}$ defined by $\tilde{\alpha}(x,y)=(\alpha_{y}x,y)$, where $\alpha\colon S^{1}\to\mathrm{SO}(n)$ represents a generator of $\pi_{1}(\mathrm{SO}(n))$. In general, the operations $\Sigma_{e}$ and $\widetilde{\Sigma}_{e}$ do not coincide. For example, if $M$ is a simply-connected non-spin manifold and $e\in H^{2}(M;\mathbb{Z})$ satisfies $e\equiv w_{2}(M)\mod 2$, then $\Sigma_{e}M$ is spin, while $\widetilde{\Sigma}_{e}M$ is non-spin, see [16, Lemma 5.2].
@@ 3.2. The metric on $[0,s_{\lambda}]\times S^{n-1}\times S^{1}$
which implies that (3.4) holds. Moreover, since $\tilde{h}^{\prime\prime}\leq 0$ and $\tilde{h}^{\prime}\in[0,h^{\prime}(s_{\lambda}-\varepsilon)]$, while $\tilde{h}(s_{\lambda})\to h(s_{\lambda})$ as $\varepsilon\to 0$ and $f^{\prime\prime}<0$ in a small neighbourhood of $s=s_{\lambda}$, also (3.2) and (3.3) hold on $[s_{\lambda}-\varepsilon,s_{\lambda}]$ for $\varepsilon$ sufficiently small.
@@
Since the functions $f,h_{r}$ satisfy inequalities (3.2)–(3.4), and since $h_{r}=r\cdot h$ on $[\varepsilon,s_{\lambda}]$, the Ricci curvatures $\textup{Ric}(T,T)$, $\textup{Ric}(X/f,X/f)$ and $\textup{Ric}(\partial_{s},\partial_{s})$ are strictly positive for $r$ sufficiently small and converge to a positive function as $r\to 0$. The mixed Ricci curvatures $\textup{Ric}(T,X/f)$, $\textup{Ric}(T,\partial_{s})$ and $\textup{Ric}(X/f,\partial_{s})$ can be bounded above by any positive constant by possibly choosing $r$ smaller. It follows that $g_{f,h_{r}}$ has positive Ricci curvature for $r$ sufficiently small. ∎
@@ 4.1. Isometric torus actions
see e.g. [16, Lemma 2.3]. Further, by [16, Corollary 2.6], the manifold $P$ is spin if and only if $w_{2}(B)\in H^{2}(B;\mathbb{Z}/2)$ is contained in the subspace generated by $e(\pi)\mod 2$. Thus, by choosing a suitable Euler class, we can realize $P$ as either a spin or non-spin manifold whenever $k>0$.
@@
The manifolds constructed in the proof of Theorem B are total spaces of principal torus bundles $P\to M_{e,\ell^{\prime}}$ over the manifolds $M_{e,\ell^{\prime}}$. It follows from [16, Theorems B and C] that each total space $P$ is diffeomorphic to a connected sums of products of spheres and possibly a non-trivial linear sphere bundle over $S^{2}$ (in case $P$ is non-spin). The existence of Riemannian metrics of positive Ricci curvature (without any assumption on symmetries) on manifolds of this form has already been established in [9] in combination with [37, Theorem C].
@@ 4.2. Connected sums and principal $S^{1}$-bundles
By [16, Theorem A], we can decompose $P$ as follows:
@@
By [16, Theorem B], each twisted suspension $\Sigma_{x_{i}}(\pm\mathbb{C}P^{2m})$ is diffeomorphic to $S^{2}\times S^{4m-1}$ and, since $x_{\ell}\in H^{2}(\mathbb{C}P^{2m})$ is a generator, we have $P_{0}\cong S^{4m+1}$. Hence $P$ is diffeomorphic to
@@
For (2) let $M^{n}$ be a closed, simply-connected manifold that admits a Riemannian metric of positive Ricci curvature and let $P\to M$ be a principal $S^{1}$-bundle with primitive Euler class $e$. Then, by [16, Theorem B], we have
@@
admits a Riemannian metric of positive Ricci curvature. The claim now follows from the fact that the manifold $N\#_{\ell-1}(S^{2}\times S^{n-1})$ is diffeomorphic to #$\#_{\ell}N$, see e.g. [16, Corollary 4.2]. ∎
@@
Note that, in order to apply the results of [16, Theorem B], we do not need that $M$ is simply-connected in (2) of Corollary C. Indeed, we only need to require that the fibre inclusion $S^{1}\hookrightarrow P$ is null-homotopic, or, equivalently, that the pull-back of the Euler class to the universal cover of $M$ is primitive. In this case we define $N$ as $S^{2}\times S^{n-1}$, whenever the universal cover of $M$ is non-spin and as the unique non-trivial linear $S^{n-1}$-bundle over $S^{2}$ otherwise.
@@ 4.3. Simply-connected 6-manifolds
$H_{2}(M_{k};\mathbb{Z})\oplus H_{2}(N_{k}\#_{7}(S^{2}\times S^{3});\mathbb{Z})\cong{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.}\oplus{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.}\oplus\mathbb{Z}^{7}.$
@@
where $e_{i}\in H^{2}(S^{2}\times S^{3};\mathbb{Z})$ denotes the restriction of $e$ to the $i$-th summand. Further, by [16, Theorem B], each $\Sigma_{e_{i}}(S^{2}\times S^{3})$ is diffeomorphic to either $(S^{2}\times S^{4})\#(S^{3}\times S^{3})$ or $(S^{2}\mathrel{\widetilde{\times}}S^{4})\#(S^{3}\times S^{3})$, where $S^{2}\mathrel{\widetilde{\times}}S^{4}$ denotes the total space of the unique non-trivial $S^{4}$-bundle over $S^{2}$, depending on whether $e_{i}$ has even or odd divisibility.
@@
Further, for any coefficient ring, the product of any two classes in $H^{2}(\Sigma_{0}N_{k})$ vanishes. This can be seen by considering the space $M^{\prime}=(S^{2}\mathrel{\widetilde{\times}}S^{3})\#N_{k}$, where $S^{2}\mathrel{\widetilde{\times}}S^{3}$ denotes the total space of the unique non-trivial linear $S^{3}$-bundle over $S^{2}$. If $x^{*}\in H^{2}(S^{2}\mathrel{\widetilde{\times}}S^{3};\mathbb{Z})\cong\mathbb{Z}$ denotes a generator, then the principal $S^{1}$-bundle with Euler class $x^{*}$ over $M^{\prime}$ is diffeomorphic to $(S^{3}\times S^{3})\#\Sigma_{0}N_{k}$ by [16, Theorem A]. By the Gysin sequence, the bundle projection induces a surjective map
@@ 4.4. Non-simply-connected rational homology spheres
$H^{i}(L_{k};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,n-1,\\ {\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&2\leq i\leq n-2\text{ and }i\text{ even},\\ 0,\quad else.\end{cases}$
@@
More generally, since any spherical space form $S_{G}=S^{n}/G$ is a rational homology sphere, the twisted suspension $\Sigma_{e}S_{G}$ is a rational homology sphere and we have $\pi_{1}(\Sigma_{e}S_{G})\cong\pi_{1}(S_{G})\cong G$ by Lemma 2.5. The universal cover of this space can be described as follows:
@@
Since $\Sigma_{0}S^{n}\cong S^{n+1}$, cf. [16, Example 5.3], we obtain that the universal cover of $\Sigma_{e}S_{G}$ is given by
@@ 4.5. Manifolds with prescribed third homology
$H^{i}(\Sigma_{e}\mathbb{C}P^{m};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2,2m-1,2m+1,\\ {\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&4\leq i\leq 2m-2,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
@@
$H^{i}(P;\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2m-1,\\ {\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&2\leq i\leq 2m-2,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
This can alternatively also be seen from the fact that $\mathbb{C}P^{m}\setminus D^{2m}$ is homotopy equivalent to $\mathbb{C}P^{m-1}$ and therefore $P$ is homotopy equivalent to a lens space of dimension $(2m-1)$ with fundamental group $\mathbb{Z}/k$. The claim now follows from Lemma 2.5.
@@
$G\cong{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{${k_{1_{\phantom{1}}}}\mathbb{Z}$}\right.}\oplus\dots\oplus{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}\right.}\oplus\mathbb{Z}^{\ell_{2}}.$
@@
$H^{i}(M_{G};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2m+1,\\ \mathbb{Z}^{\ell},\quad&i=2,2m-1,\\ \mathbb{Z}^{\ell_{2}},\quad&i=3,\\ G,\quad&i=2m-2,\\ {\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{${k_{1_{\phantom{1}}}}\mathbb{Z}$}\right.}\oplus\dots\oplus{\raisebox{1.99997pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}\right.},\quad&4\leq i\leq 2m-4,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
@@ Appendix A Cohomogeneity-two torus actions
see e.g. [16, Lemma 2.3].
Furthermore, we have $b_{2}(M^{\prime})=0$, since otherwise, again by [16, Lemma 2.3], we can consider a principal $S^{1}$-bundle over $M^{\prime}$ with simply-connected total space $P^{\prime}$. By lifting the action from $M^{\prime}$ to $P^{\prime}$, see [21, Proposition 6.2], we obtain an effective action of $T^{m+1}$ on $P^{\prime}$ with a free sub-action of a circle. Since the action is free, dividing out this action does not change the number of vertices in the orbit space. Thus, the orbit space of $P^{\prime}$ is an $m$-gon labelled by $m$-elements of $\mathbb{Z}^{m+1}$ that generate $\mathbb{Z}^{m+1}$, which is a contradiction.

### v3

@@ 1. Introduction and Main Results
$G\cong{\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{1}\mathbb{Z}$}}\oplus\dots\oplus{\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}}\oplus\mathbb{Z}^{\ell_{2}}$
@@ Acknowledgements.
The author would like to thank Fernando Galaz-García and Philippe Kupper for helpful discussions, and them as well as Sam Hagh Shenas Noshari, Michael Wiemeler and the anonymous referee for helpful comments on an earlier version of this article. The author would also like to thank the Department of Mathematical Sciences of Durham University for its hospitality during a stay where parts of this work were carried out.
@@ 2.2. Twisted suspensions
In [18] another suspension operation, denoted $\widetilde{\Sigma}_{e}$, is introduced, which is obtained by replacing the identity map on $S^{n-1}\times S^{1}$ in Definition 2.4 by a map $\tilde{\alpha}\colon S^{n-1}\times S^{1}\to S^{n-1}\times S^{1}$ defined by $\tilde{\alpha}(x,y)=(\alpha_{y}x,y)$, where $\alpha\colon S^{1}\to\mathrm{SO}(n)$ represents a generator of $\pi_{1}(\mathrm{SO}(n))$. In general, the operations $\Sigma_{e}$ and $\widetilde{\Sigma}_{e}$ do not coincide. For example, if $M$ is a simply-connected non-spin manifold and $e\in H^{2}(M;\mathbb{Z})$ satisfies $e\equiv w_{2}(M)\mod 2$, then $\Sigma_{e}M$ is spin, while $\widetilde{\Sigma}_{e}M$ is non-spin, see [18, Lemma 5.2].
@@ 3.2. The metric on $[0,s_{\lambda}]\times S^{n-1}\times S^{1}$
which implies that (3.4) holds. Moreover, since $\tilde{h}^{\prime\prime}\leq 0$ and $\tilde{h}^{\prime}\in[0,h^{\prime}(s_{\lambda}-\varepsilon)]$, while $\tilde{h}(s_{\lambda})\to h(s_{\lambda})$ as $\varepsilon\to 0$ and $f^{\prime\prime}<0$ in a small neighbourhood of $s=s_{\lambda}$, also (3.2) and (3.3) hold on $[s_{\lambda}-\varepsilon,s_{\lambda}]$ for $\varepsilon$ sufficiently small. ∎
@@
Since the functions $f,h_{r}$ satisfy inequalities (3.2)–(3.4), and since $h_{r}=r\cdot h$ on $[\varepsilon,s_{\lambda}]$, the Ricci curvatures $\textup{Ric}(T,T)$, $\textup{Ric}(X/f,X/f)$ and $\textup{Ric}(\partial_{s},\partial_{s})$ are strictly positive for $r$ sufficiently small and converge to a positive function as $r\to 0$. The absolute values of the mixed Ricci curvatures $\textup{Ric}(T,X/f)$, $\textup{Ric}(T,\partial_{s})$ and $\textup{Ric}(X/f,\partial_{s})$ can be bounded above by any positive constant by possibly choosing $r$ smaller. It follows that $g_{f,h_{r}}$ has positive Ricci curvature for $r$ sufficiently small. ∎
@@ 4.1. Isometric torus actions
see e.g. [18, Lemma 2.3]. Further, by [18, Corollary 2.6], the manifold $P$ is spin if and only if $w_{2}(B)\in H^{2}(B;\mathbb{Z}/2)$ is contained in the subspace generated by $e(\pi)\mod 2$. Thus, by choosing a suitable Euler class, we can realize $P$ as either a spin or non-spin manifold whenever $k>0$.
@@
The manifolds constructed in the proof of Theorem B are total spaces of principal torus bundles $P\to M_{e,\ell^{\prime}}$ over the manifolds $M_{e,\ell^{\prime}}$. It follows from [18, Theorems B and C] that each total space $P$ is diffeomorphic to a connected sum of products of spheres and possibly a non-trivial linear sphere bundle over $S^{2}$ (in case $P$ is non-spin). The existence of Riemannian metrics of positive Ricci curvature (without any assumption on symmetries) on manifolds of this form has already been established in [9] in combination with [37, Theorem C].
@@ 4.2. Connected sums and principal $S^{1}$-bundles
By [18, Theorem A], we can decompose $P$ as follows:
@@
By [18, Theorem B], each twisted suspension $\Sigma_{x_{i}}(\pm\mathbb{C}P^{2m})$ is diffeomorphic to $S^{2}\times S^{4m-1}$ and, since $x_{\ell}\in H^{2}(\mathbb{C}P^{2m})$ is a generator, we have $P_{0}\cong S^{4m+1}$. Hence $P$ is diffeomorphic to
@@
For (2) let $M^{n}$ be a closed, simply-connected manifold that admits a Riemannian metric of positive Ricci curvature and let $P\to M$ be a principal $S^{1}$-bundle with primitive Euler class $e$. Then, by [18, Theorem B], we have
@@
admits a Riemannian metric of positive Ricci curvature. The claim now follows from the fact that the manifold $N\#_{\ell-1}(S^{2}\times S^{n-1})$ is diffeomorphic to #$\#_{\ell}N$, see e.g. [18, Corollary 4.2]. ∎
@@
Note that, in order to apply the results of [18, Theorem B], we do not need that $M$ is simply-connected in (2) of Corollary C. Indeed, we only need to require that the fibre inclusion $S^{1}\hookrightarrow P$ is null-homotopic, or, equivalently, that the pull-back of the Euler class to the universal cover of $M$ is primitive. In this case we define $N$ as $S^{2}\times S^{n-1}$, whenever the universal cover of $M$ is non-spin and as the unique non-trivial linear $S^{n-1}$-bundle over $S^{2}$ otherwise.
@@ 4.3. Simply-connected 6-manifolds
$H_{2}(M_{k};\mathbb{Z})\oplus H_{2}(N_{k}\#_{7}(S^{2}\times S^{3});\mathbb{Z})\cong{\raisebox{0.0pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.}\oplus{\raisebox{0.0pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.}\oplus\mathbb{Z}^{7}.$
@@
where $e_{i}\in H^{2}(S^{2}\times S^{3};\mathbb{Z})$ denotes the restriction of $e$ to the $i$-th summand. Further, by [18, Theorem B], each $\Sigma_{e_{i}}(S^{2}\times S^{3})$ is diffeomorphic to either $(S^{2}\times S^{4})\#(S^{3}\times S^{3})$ or $(S^{2}\mathrel{\widetilde{\times}}S^{4})\#(S^{3}\times S^{3})$, where $S^{2}\mathrel{\widetilde{\times}}S^{4}$ denotes the total space of the unique non-trivial $S^{4}$-bundle over $S^{2}$, depending on whether $e_{i}$ has even or odd divisibility.
@@
Further, for any coefficient ring, the product of any two classes in $H^{2}(\Sigma_{0}N_{k})$ vanishes. This can be seen by considering the space $M^{\prime}=(S^{2}\mathrel{\widetilde{\times}}S^{3})\#N_{k}$, where $S^{2}\mathrel{\widetilde{\times}}S^{3}$ denotes the total space of the unique non-trivial linear $S^{3}$-bundle over $S^{2}$. If $x^{*}\in H^{2}(S^{2}\mathrel{\widetilde{\times}}S^{3};\mathbb{Z})\cong\mathbb{Z}$ denotes a generator, then the principal $S^{1}$-bundle with Euler class $x^{*}$ over $M^{\prime}$ is diffeomorphic to $(S^{3}\times S^{3})\#\Sigma_{0}N_{k}$ by [18, Theorem A]. By the Gysin sequence, the bundle projection induces a surjective map
@@ 4.4. Non-simply-connected rational homology spheres
$H^{i}(L_{k};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,n-1,\\ {\raisebox{0.0pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&2\leq i\leq n-2\text{ and }i\text{ even},\\ 0,\quad else.\end{cases}$
@@
More generally, since any orientable spherical space form $S_{G}=S^{n}/G$ is a rational homology sphere, the twisted suspension $\Sigma_{e}S_{G}$ is a rational homology sphere and we have $\pi_{1}(\Sigma_{e}S_{G})\cong\pi_{1}(S_{G})\cong G$ by Lemma 2.5. The universal cover of this space can be described as follows:
@@
Since $\Sigma_{0}S^{n}\cong S^{n+1}$, cf. [18, Example 5.3], we obtain that the universal cover of $\Sigma_{e}S_{G}$ is given by
@@ 4.5. Manifolds with prescribed third homology
$H^{i}(\Sigma_{e}\mathbb{C}P^{m};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2,2m-1,2m+1,\\ {\raisebox{0.0pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&4\leq i\leq 2m-2,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
@@
$H^{i}(P;\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2m-1,\\ {\raisebox{0.0pt}{$\mathbb{Z}$}\left/\raisebox{-1.99997pt}{$k\mathbb{Z}$}\right.},\quad&2\leq i\leq 2m-2,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
This can alternatively also be seen from the fact that $\mathbb{C}P^{m}\setminus D^{2m}$ is homotopy equivalent to $\mathbb{C}P^{m-1}$ and therefore $P$ is homotopy equivalent to a lens space of dimension $(2m-1)$ with fundamental group $\mathbb{Z}/k$. The claim now follows from Lemma 2.5. ∎
@@
$G\cong{\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{1}\mathbb{Z}$}}\oplus\dots\oplus{\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}}\oplus\mathbb{Z}^{\ell_{2}}.$
@@
$H^{i}(M_{G};\mathbb{Z})\cong\begin{cases}\mathbb{Z},\quad&i=0,2m+1,\\ \mathbb{Z}^{\ell},\quad&i=2,2m-1,\\ \mathbb{Z}^{\ell_{2}},\quad&i=3,\\ G,\quad&i=2m-2,\\ {\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{1}\mathbb{Z}$}}\oplus\dots\oplus{\raisebox{0.0pt}{$\mathbb{Z}$}\big{/}\raisebox{-1.99997pt}{$k_{\ell_{1}}\mathbb{Z}$}},\quad&4\leq i\leq 2m-4,\text{ and }i\text{ even},\\ 0,\quad&\text{else.}\end{cases}$
@@ Appendix A Cohomogeneity-two torus actions
see e.g. [18, Lemma 2.3].
Furthermore, we have $b_{2}(M^{\prime})=0$, since otherwise, again by [18, Lemma 2.3], we can consider a principal $S^{1}$-bundle over $M^{\prime}$ with simply-connected total space $P^{\prime}$. By lifting the action from $M^{\prime}$ to $P^{\prime}$, see [21, Proposition 6.2], we obtain an effective action of $T^{m+1}$ on $P^{\prime}$ with a free sub-action of a circle. Since the action is free, dividing out this action does not change the number of vertices in the orbit space. Thus, the orbit space of $P^{\prime}$ is an $m$-gon labelled by $m$-elements of $\mathbb{Z}^{m+1}$ that generate $\mathbb{Z}^{m+1}$, which is a contradiction.